OpenStudy (anonymous):

5/(x-4) + 3/(x+4) = -6x/(x^2-16) solve for x

OpenStudy (vijay):

-0.5714

OpenStudy (anonymous):

factor first. x^2-16 is the difference of 2 perfect squares, so (x-4)(x+4)

OpenStudy (anonymous):

. vijay that says wrong

OpenStudy (vijay):

cannot be

OpenStudy (anonymous):

If u multiply the whole equation by (x-4)(x+4), all ur denominators go away

OpenStudy (anonymous):

that's what it says

OpenStudy (anonymous):

maybe no negative?

OpenStudy (vijay):

x = -4/ 7

OpenStudy (anonymous):

5(x+4)+3(x-4)=6x

OpenStudy (vijay):

-4/7 = -0.5714

OpenStudy (anonymous):

5x+20+3x-12=6x

OpenStudy (anonymous):

it is -4/7

OpenStudy (anonymous):

Move the RHS to the LHS\[\frac{6 x}{x^2-16}+\frac{5}{x-4}+\frac{3}{x+4}=0 \]Combine fractions and simplify.\[\frac{2 (7 x+4)}{(x-4) (x+4)}=0 \]